OpenStudy (codingking03):

Solve for x: 5x + 2 = 4x − 9.

OpenStudy (giantrobot):

This is algebra. Always remember you can move values. Like so. \[A=B\] \[A-B=B-B\] is what your're doing when you move a value in algebra. Lets apply this to your Question. \[5x+2=4x-9\] \[5x+2-4x=4x-9-4x\] \[5x+2-4x-2=4x-9-4x-2\] Work that equation.

OpenStudy (codingking03):

thx